OpenStudy (anonymous):

How can I show that \(\phi(n)=14\) is impossible? Here, \(\phi\) is Euler's phi-function.

hero (hero):

pre-algebra, I'm never helping you with any of your questions.

OpenStudy (anonymous):

Why not? :(

hero (hero):

Because, everytime you post a question, it is something I've never seen before.

hero (hero):

What course are you taking?

OpenStudy (anonymous):

I have an idea for this one: Suppose that there is a number \(n\) such that \(\phi(n)=14\). Then \(n\) has prime-power factorization \(n=p_{1}^{a_{1}}p_{2}^{a_{2}}\cdots p_{k}^{a_{k}}\) and \(\phi(n)=\phi(p_{1}^{a_{1}}p_{2}^{a_{2}}\cdots p_{k}^{a_{k}})=\phi(p_{1}^{a_{1}})\phi(p_{2}^{a_{2}})\cdots \phi(p_{k}^{a_{k}})=14\). Moreover, the prime-power factorization of \(14\) is \(14=2\cdot7\). We then need (by definition) two numbers of the form \(p^a(p-1)\) with \(p\) prime and \(a\) a positive integer. Clearly, \(2=2^1(2-1)\), but 7=p^a(p-1) would imply that \(7\) is composite since \(2\) is the smallest prime such that \(2-1=1\) and \(2^3=8>7\). This is elementary number theory.

hero (hero):

That would explain why I could never help you.